We call a positive integer perfect if it equals the sum of its positive divisors


other than itself.


(a) Prove that 6 and 28 are perfect numbers


(b) Prove that if 2p − 1 is prime, then 2p−1



1
Expert's answer
2021-11-10T17:44:47-0500

a.

The divisors of "6" are "1,2, \space and \space 3". If we sum these divisors then we have ,

"1+2+3=6", Therefore, 6 is a perfect number.

The divisors of "28" are "1,2,4,7,14". Summing these divisors gives,

"1+2+4+7+14=28". This shows that "28" is a perfect number.


b.

We need to show that If "2^p-1" is a prime number, then "2^{p-1}(2^p-1)" is a perfect number. 

Let "p" be an integer so that "2^p-1" is a prime number. Also, let "n=2^p-1" so that "2^{p-1}(2^p-1)=2^{p-1}n". We can write the divisors of "2^{p-1}" as, "1,2,4,8,...,2^{p-1}". Since "n"

is prime, we can write out the other divisors of "2^{p-1}n" as, "n,2n,4n,8n,...,2^{p-2}n".

Let us recall that,

"1+x+x^2+x^3+...+x^{p-1}=(x^p-1)\/(x-1)", where "x=2" so,

"1+2+4+8+...+2^{p-1}=(2^p-1)\/(2-1)=2^p-1=n...........(i)"

We can use the same formula we have used to find "equation(i)" for,

"n+2n+4n+8n+...+2^{p-2}n"

We can see that,

"n+2n+4n+8n+...+2^{p-2}n=n(1+2+4+8+...+2^{p-2})=n(1+2^1+2^2+2^3+...+2^{p-2})=n(2^{p-1}-1)\/(2-1)=n(2^{p-1}-1)............(ii)"

We can now sum all the divisors of "2^{p-1}n" by combining equations "(i)" and "(ii)", "1+2+4+...+2^{n-1}+n+2n+4n+...+2^{p-2}n=n+n(2^{p-1}-1)=n+2^{p-1}n-n=2^{p-1}n=2^{p-1}(2^p-1)"

Therefore, if "2^p-1" is prime then "2^{p-1}(2^p-1)" is perfect. "\blacksquare"